OpenStudy (anonymous):

A boat moves 9 kilometers upstream in the same amount of time it moves 19 kilometers downstream. If the rate of the current is 9 kilometers per hour, find the rate of the boat in still water.

OpenStudy (anonymous):

\[\frac{9}{r-9} = \frac{19}{r+9}\] ?

OpenStudy (anonymous):

These are the Answer choices! A. 17 km/hr B. 9 km/hr C. 8 km/hr D. 25 km/hr

OpenStudy (anonymous):

9r + 81 = 19r - 171 10r = 252 r = 25.2 Answer should be either A or D, but I think it's D

OpenStudy (anonymous):

Oh I have no ideal! LOL

OpenStudy (anonymous):

Correct answer is D.

OpenStudy (anonymous):

So my equation was right

OpenStudy (anonymous):

Yes! Thank you!